OpenStudy (anonymous):

can someone help me show this: lim (((n+1)^m )-n^m) n->+inf ---------------- = m n^(m-1)

OpenStudy (anonymous):

sorry, i don't understand the front part of: x--+oo

OpenStudy (anonymous):

\[\LARGE \lim_{x \rightarrow +\infty} \frac{(((n+1)^m )-n^m) }{n^(m-1)}\]

OpenStudy (anonymous):

but the point is where is X in the question?

OpenStudy (anonymous):

The denominator wnats in the parantheses is in exposant

OpenStudy (anonymous):

No its show that this equation = m

OpenStudy (anonymous):

\[\LARGE \lim_{x \rightarrow +\infty} \frac{(((n+1)^m )-n^m) }{n^{(m-1)}}\]

OpenStudy (anonymous):

limit x tending to infinity,the question is wrong

OpenStudy (anonymous):

Sorry its n

OpenStudy (anonymous):

\[\LARGE \lim_{n \rightarrow +\infty} \frac{(((n+1)^m )-n^m) }{n^{(m-1)}}\]

OpenStudy (anonymous):

thats right .

OpenStudy (anonymous):

Do you know how it equals m ?